# Nexus: Taming Throughput-Latency Tradeoff in LLM Serving via Efficient GPU Sharing

### v2

@@ 3.2. Diminishing Returns in Compute Allocation
While the Section 3.1 highlights the limitations of both chunked prefill and inter-engine PD disaggregation, we take a step further by exploring an intra-engine PD approach through GPU resource sharing. In this design, prefill and decode are assigned different portions of the same GPU. To allocate GPU resources effectively under prefill–decode separation, we analyze how each phase scales with compute in isolation. We run pure prefill and decode batches under varying SM ratios, measuring both end-to-end latency and per-kernel runtimes.
As shown in Figure 5, prefill latency closely follows the idealized scaling model $T\propto\frac{1}{r}$, with diminishing returns emerging gradually. For instance, increasing SM allocation from 30% to 40% reduces latency by over 25%, but the gain drops to just 10% between 70% and 80%. Further investigation in Figure 5(b) shows that compute-heavy layers such as FFN continue to benefit, while others such as KQV and projection flatten out earlier, around 60%.
Decode exhibits much sharper diminishing returns. Increasing SMs from 30% to 40% improves latency by only 10%, and beyond 50%, additional SMs yield less than 3% improvement per 10% increment. This behavior is expected given decode is memory-bound and fails to utilize added SMs effectively. Figure 5(c) confirms this: attention and projection layers show minimal runtime reduction with more compute.
These trends suggest a key systems insight: instead of time-slicing all SMs between the two stages, we can spatially partition them to avoid overprovisioning either phase. Notably, disaggregated systems, which allocate entire GPUs to each phase, operate on the far right of these curves, where additional compute offers diminishing marginal benefit.
@@
[backgroundcolor=gray!7, linewidth=0.8pt] Insight 2. Both prefill and decode saturate well before full GPU allocation. To improve efficiency, systems should allocate only the SMs needed to meet each phase’s demand.
@@ 3.3. Limitation of Static Partitioning
Section 3.2 shows that prefill and decode phases exhibit sharply diminishing returns with increasing SM allocation, motivating finer-grained partitioning. However, even a well-chosen static SM partition is suboptimal at runtime. This is because compute demand, while often estimable from inputs like chunk size or sequence length, does not capture dynamic memory behavior. Prefill and decode contend for bandwidth in ways that depend on evolving KV cache sizes, prompt distributions, and decode lengths that emerge only at runtime.
To illustrate this, we co-execute chunks of prefill [^2] Following Sarathi-Serve (Agrawal et al., 2024), long prompts are split into smaller segments and executed alongside decode requests. These chunks of prefill tasks read and write to the KV cache. and pure decodes under a fixed SM partition. As we can see from Figure 6(a), increasing the prefill KV length from 2000 to 10000 also increases the latency of the exact same decode batch by 36%. This slowdown stems from memory bandwidth contention: prefill attention layers perform large KV reads, which overlap with the decode stage’s latency-critical memory access.
Moreover, Figure 6(b) shows that prefill memory traffic is highly irregular, fluctuating significantly over time and making contention both difficult to predict and workload dependent.
These highlight a core limitation: even when compute demands are static and analyzable, bandwidth pressure varies at runtime due to asynchronous prefill/decode evolution. Static SM partitioning fails to respond to this variation, leading to avoidable contention and degraded latency.
@@
[backgroundcolor=gray!7, linewidth=0.8pt] Insight 3. Logical PD disaggregation with static partitioning is insufficient. To adapt to emergent memory contention and shifting runtime demands, systems must have fine-grained, dynamic SM reallocation.
@@ 4. Design
We present Nexus, a lightweight monolithic LLM serving system that enables intra-engine prefill–decode disaggregation. Unlike prior systems that either co-batch prefill and decode or isolate them across GPUs, Nexus partitions GPU resources dynamically, executing both phases concurrently but independently within a single engine.
As shown in Figure 7, Nexus introduces three core mechanisms to enable this fine-grained separation:
@@ 4.1. Dynamic SM Partitioning via Cost Model
Nexus’s cost model estimates the latency of prefill and decode under SM allocations without execution, enabling rapid exploration of latency tradeoffs.
@@
Compute Latency. We estimate the compute latency of each operator $o\in\texttt{PrefillOps}\cup\texttt{DecodeOps}$ based on its FLOP count $c_{o}$ and the SM ratio $r$ assigned to its stage. While latency ideally scales inversely with compute share ($1/r$), real-world performance can deviates depending on kernel(§3.2).
@@
We extract $R_{\text{sat}}$ and $\lambda$ per operator from end-to-end measurements of the full stage (prefill or decode) under varying SM allocations.
Memory Access Latency. As shown in §3.3, memory contention can significantly affect latency when prefill and decode execute concurrently on shared hardware. To capture this effect, we model decode memory latency as a function of (1) temporal overlap with prefill, and (2) the relative memory bandwidth demands of each phase.
@@
Assuming full bandwidth saturation during each overlap window, we allocate effective bandwidth to decode based on its share of memory traffic. Let $m_{d}$ be the total memory bytes accessed by decode attention, $m_{p1}$ the bytes accessed by prefill attention, and $m_{p2}$ those accessed by prefill’s dense operators. [^3] We estimate memory access volume for linear layers using known parameter sizes from the model architecture. For attention, we calculate KV memory traffic based on the number of cached tokens tracked by vLLM, multiplied by the per-token key/value size determined from model dimensions. Then the effective bandwidth for decode is:
@@
Given the cost model, Nexus selects an SM partition that balances performance and memory pressure. However, since prefill and decode run concurrently and compete for resources, optimizing both simultaneously is infeasible.
To resolve this, Nexus formulates a dual-objective latency optimization: prioritize one phase while constraining the other to remain within a slowdown budget. This allows flexible tradeoffs between TBT and TTFT, depending on workload state. The choice is also guided by runtime signals, such as KV cache usage, to avoid pathological behaviors like OOM.
@@
Here, $T^{\text{min}}$ denotes the ideal latency when a stage is allocated all SMs, and $\alpha,\beta>1$ are slack variables controlling tolerable slowdowns in the non-prioritized stage.
Runtime Switching. We select the objective mode based on live KV cache usage $KV_{u}$: when $KV_{u}$ is low, favoring prefill accelerates prompt ingestion; when $KV_{u}$ is high, prioritizing decode helps reduce memory pressure by completing generations and evicting KV. This feedback mechanism enables resource allocation to respond to workload phases and memory constraints.
@@ 4.1.3. Greedy SM Search
Efficiency. The search typically converges within 2–4 cost model evaluations and imposes negligible latency overhead. This fast feedback loop allows Nexus to adapt to runtime contention and shifting workloads.
@@ 4.2. Hysteresis-Based SM Repartitioning for Stability
This simple mechanism retains the adaptability of fine-grained decisions while avoiding the instability of overreactive switching. Alternative approaches like reducing update frequency or increasing step size proved brittle in practice: the former sacrifices responsiveness; the latter risks overshooting optimal ratios. Buffered switching provides a robust, low-overhead tradeoff.
@@ 4.3. Phase-Specific Scheduler
Since Nexus separates prefill and decode into concurrent batches, we can further exploit this with phase-specific optimizations. Particularily, we employee tailored scheduler policies that address the distinct latency and resource profiles.
@@ 4.3.2. Decode Scheduler
The decode phase controls TBT. Unlike prefill, decode scheduling operates at a finer granularity, with each request contributing only a single token to the active batch. Thus, we adopt a simple First-Come-First-Serve (FCFS) policy. FCFS ensures fairness, incurs minimal overhead, and avoids token-level starvation. While more sophisticated schemes that considers context window size or memory bandwidth are theoretically possible, we find that their impact is limited in practice as they are considered in SM partitioning.
@@ 6.1. Experimental Setup
Baselines. We compare Nexus against four representative LLM serving engines, all configured with the same tensor parallelism, chunked prefill, and scheduling budget for fair comparison.
@@ 7. Related Work
Monolithic LLM Serving Systems. Early LLM serving engines focus on maximizing throughput and memory efficiency within a unified execution model. Orca (Yu et al., 2022) introduces continuous batching to reduce head-of-line blocking. vLLM (Kwon et al., 2023) eliminates KV cache fragmentation via PagedAttention, while SGLang (Zheng et al., 2023) reduces memory usage through RadixAttention. SarathiServe (Agrawal et al., 2024) mixes chunks of prefill and decode requests to better utilize GPU resources. However, all of these designs treat prefill and decode as indistinguishable units within a shared queue. In contrast, Nexus decouples the two phases at the batching level, enabling independent, phase-specific execution and scheduling while preserving compatibility with existing attention mechanisms.
@@
Disaggregated Serving Architectures. Several systems physically disaggregate prefill and decode across GPUs to better match their resource profiles. Splitwise (Patel et al., 2024) statically assigns phases to different hardware tiers; DistServe (Zhong et al., 2024) improves TTFT/TBT via tiered GPU scheduling; Mooncake (Qin et al., 2025) serves cached KV blocks from storage to reduce compute load; TetriInfer (Hu et al., 2024) routes requests by latency class to isolated replicas. Nexus achieves similar benefits without incurring cross-engine complexity. By performing lightweight intra-engine disaggregation, it spatially partitions resources for concurrent phase-aware execution—enabling low-latency decode and efficient KV reuse within a single serving engine.
LLM Scheduling Frameworks. Recent works propose sophisticated schedulers to balance latency and throughput. FastServe (Wu et al., 2023) uses MLFQ with skip-join to avoid prompt variance stalls. VTC (Sheng et al., 2024) and QLM (Patke et al., 2024) target fairness and SLO adherence. Llumnix (Sun et al., 2024) accelerates dynamic scaling via migration-aware scheduling. LightLLM (Gong et al., 2025) and Preble (Srivatsa et al., 2024) improve memory reuse via future prediction or prompt sharing. Yet, all treat each request as a monolithic scheduling unit. Nexus introduces dual schedulers for prefill and decode, each tailored to its phase’s latency sensitivity and compute intensity. This decoupling enables tighter queue control and more efficient GPU resource utilization.
Intra-Engine PD Disaggregation. There are some prior works on intra-engine PD disaggregation (Cui et al., 2025; Lin et al., 2025). Drift (Cui et al., 2025) introduces an adaptive gang scheduling mechanism, a contention-free performance model, and an SLO-aware dispatching policy to enable intra-engine prefill–decode separation. Bullet (Lin et al., 2025) proposes a comprehensive system that includes: (1) a performance estimator for building a profile-augmented analytical model; (2) an SLO-aware scheduler to dynamically balance the compute load between prefill and decode phases; and (3) a resource manager capable of delivering fast yet accurate resource configurations. The key differences between their and Nexus are as follows: Nexus employs a contention-based cost model to estimate latency, formulates a dual-objective optimization problem, and uses a greedy search to determine the optimal SM partitioning. In addition, our scheduler is phase-aware and explicitly considers the distinct characteristics of prefill and decode stages, which contrasts with Drift’ SLO-aware scheduling approach.
@@
GPU Multiplexing. Recent work explores fine-grained GPU sharing via spatial or temporal partitioning. NVIDIA MPS (Nvidia, 2025e) and MIG (Nvidia, 2025d) provide coarse-grained isolation, while GreenContext (Nvidia, 2025b) enables dynamic intra-process SM partitioning. Systems like GPUlet, Orion, REEF, and Bless (Choi et al., 2022; Strati et al., 2024; Han et al., 2022; Zhang et al., 2025) improve utilization for small models through temporal multiplexing. MuxServe (Duan et al., 2024) proposes spatial-temporal multiplexing to efficiently serve multiple LLMs, while NanoFlow (Zhu et al., 2024) and Liger (Du et al., 2024) introduce kernel level parallelism that enables concurrent execution of compute bound, network bound, and memory bound operations. PoD (Kamath et al., 2025) fuses prefill attention and decode attention into one kernel to reduce overhead.
Nexus differs in three key aspects: (1)it disaggregates prefill and decode into different batch and executes these batch concurrently; and (2) it dynamically reallocate SMs across prefill and decode stages in a single engine, adapting to workload shifts; and (3) it introduces intra-engine, phase-specific schedulers for coordinated but decoupled execution. This design enables fine-grained responsiveness while maximizing intra-GPU parallelism.
@@ 8. Conclusion
While the asymmetric resource demands of prefill and decode phases in LLM inference is well recognized, its implications for intra-GPU coordination have remained underexplored. In this work, we analyze how co-executing these phases under coarse-grained batching strategies leads to resource imbalance and performance interference. To address this, we propose Nexus, a novel SM-partitioned execution framework with dynamic resource allocation and phase-specific schedulers. Our system consistently improves end-to-end throughput and latency across diverse LLM workloads, outperforming state-of-the-art serving systems.

### v3

@@ 3.2. Diminishing Returns in Compute Allocation
To guide effective GPU resource allocation between prefill and decode, we begin by analyzing how each phase scales with compute in isolation. We run pure prefill and decode batches under varying SM ratios, measuring both end-to-end latency and per-kernel runtimes.
As shown in Figure 5, prefill latency follows the expected scaling curve $T\propto\frac{1}{r}$, with diminishing returns emerging gradually. For example, increasing SM allocation from 30% to 40% reduces latency by over 25%, but this gain drops to 10% when increasing from 70% to 80%. A breakdown in Figure 5(b) shows that compute-heavy layers such as FFN continue to benefit from more SMs, while others like KQV and projection flatten earlier at 60%.
Decode exhibits sharper diminishing returns. Increasing SMs from 30% to 40% yields only a 10% latency reduction, and beyond 50%, further increases offer less than 3% improvement per 10% step. This behavior aligns with the decode phase being memory-bound and less compute-scalable. As shown in Figure 5(c), attention and projection kernels see minimal speedup.
These trends highlight a critical asymmetry: prefill and decode have different compute saturation points. Instead of time-slicing all SMs between the two stages, it is more efficient to spatially partition resources, assigning only what each phase needs. Engine-Level Disaggregated systems, which dedicate full GPUs to each phase, often operate in high regions of diminishing returns.
@@
[backgroundcolor=gray!7, linewidth=0.8pt] Insight 1. Both prefill and decode saturate well before full GPU allocation. To avoid overprovisioning, systems should allocate only the SMs needed to meet each phase’s demand. These compute asymmetries motivate analytical modeling that accounts for per-kernel scaling to guide partitioning.
@@ 3.3. Limitation of Static Partitioning
Section 3.2 shows that prefill and decode exhibit diminishing returns with increasing SM allocation, motivating phase-aware compute partitioning. However, even a well-chosen static partition is insufficient in practice. This is because compute demand, while often estimable from inputs like chunk size or sequence length, does not account for dynamic memory behavior during serving. Prefill and decode phases contend for memory bandwidth in ways that depend on runtime factors such as evolving KV-cache sizes, prompt distributions, and decode lengths.
To illustrate this, we co-execute prefill chunks [^2] Following Sarathi-Serve (Agrawal et al., 2024), long prompts are split into smaller segments and processed alongside decode requests. These prefill chunks actively read and write to the KV cache. with pure decode batches under a fixed SM partition. As shown in Figure 6(a), increasing the prefill KV length from 2,000 to 10,000 increases decode latency by 36%, despite the decode workload itself remaining constant. This slowdown stems from bandwidth contention: prefill attention layers issue large KV reads that overlap with latency-critical memory accesses from the decode phase.
Figure 6(b) further shows that prefill memory traffic is highly irregular and fluctuates significantly over time, making contention both unpredictable and workload-dependent.
These results highlight a fundamental limitation: even when compute needs are analyzable, memory bandwidth pressure can vary rapidly due to asynchronous phase behavior. Static SM partitioning cannot react to this dynamic interference, leading to resource underutilization and inflated latency.
@@
[backgroundcolor=gray!7, linewidth=0.8pt] Insight 2. Intra-GPU PD disaggregation with static resource allocation is insufficient under dynamic memory contention. Effective intra-GPU scheduling requires runtime-aware, fine-grained SM reallocation that adapts to emergent phase interference.
@@ 4. Design
We present Nexus, a lightweight monolithic LLM serving system that enables intra-GPU prefill–decode disaggregation through proactive SM partitioning and phase-aware scheduling. Unlike reactive systems that adjust only after SLO violations, Nexus proactively predicts performance bottlenecks using a contention-aware cost model and dynamically reallocates resources in real time. This enables high utilization without introducing interference. As shown in Figure 7, Nexus introduces three core mechanisms to enable this fine-grained separation:
@@ 4.1. Dynamic SM Partitioning via Cost Model
Nexus’s cost model estimates the latency of prefill and decode under SM allocations without execution. This avoids reliance on feedback-based tuning or runtime latency profiling, and supports proactive, per-batch optimization not possible in prior reactive systems.
@@
Compute Latency. We estimate the compute latency of each operator $o\in\texttt{PrefillOps}\cup\texttt{DecodeOps}$ based on its FLOP count $c_{o}$ and the SM ratio $r$ assigned to its stage. While latency ideally scales inversely with compute share ($1/r$), real-world performance can deviate depending on the kernel(§3.2).
@@
These coefficients ($R_{\text{sat}},\lambda$) are extracted from a one-time profiling pass per model, avoiding the need for runtime latency fitting and transferable across workloads.
Memory Contention Aware Model. As shown in §3.3, memory contention can significantly affect latency when prefill and decode execute concurrently on shared hardware. Unlike prior work that ignores phase overlap or assumes fixed latency models, we model decode attention latency as a function of runtime overlap with memory-bound prefill activity and the relative memory bandwidth during contention.
@@
Assuming full bandwidth saturation during each overlap window, we allocate effective bandwidth to decode based on its share of memory traffic. Let $m_{d}$ be the total memory bytes accessed by decode attention, $m_{p1}$ the bytes accessed by prefill attention, and $m_{p2}$ those accessed by prefill’s dense operators. [^3] We estimate memory access volume for linear layers using known parameter sizes from the model architecture. For attention, we weigth memory access and calculate KV memory traffic based on the number of cached tokens tracked by vLLM, multiplied by the per-token key/value size determined from model dimensions. Then the effective bandwidth for decode is:
@@
Given the cost model, Nexus selects an SM partition that balances performance and memory pressure. However, since prefill and decode run concurrently and compete for GPU compute, GPU memory and GPU memory bandwidth, optimizing both simultaneously can lead to inefficient tradeoffs.
To resolve this, Nexus formulates a dual-objective latency optimization: prioritize one phase while constraining the other to remain within a slowdown budget. This allows flexible tradeoffs between TBT and TTFT, depending on workload state. The choice is also guided by runtime signals, such as KV cache usage, to avoid pathological behaviors like OOM or long waiting time.
@@
Here, $T^{\text{min}}$ denotes the ideal latency when a stage is allocated all SMs, and $\alpha,\beta>1$ are slack variables controlling tolerable slowdowns in the non-prioritized stage. Unlike SLO thresholds, these slack constraints are batch-dependent and derived from the cost model itself.
Runtime Switching. We select the objective mode based on live KV cache usage $KV_{u}$: when $KV_{u}$ is low, prioritizing prefill improves TTFT by accelerating prompt ingestion. When $KV_{u}$ is high, prioritizing decode helps reduce memory pressure by completing generation and evicting old KV entries. This feedback mechanism enables resource allocation to respond to workload phases and memory constraints.
@@ 4.1.3. Greedy SM Search
Efficiency. We avoid global constrained solvers due to their overhead. Instead, a simple greedy search provides robustness and transparency while converging in just 2–4 iterations. This efficiency is also critical for per-batch decision making during inference.
@@ 4.2. Hysteresis-Based SM Repartitioning for Stability
This simple mechanism retains the adaptability of fine-grained decisions while avoiding the instability of overreactive switching. Alternative approaches like naively reducing the update frequency sacrifices responsiveness, while large step sizes risk overshooting optimal ratios. In contrast, our hysteresis design filters out noise and enables smooth SM reallocation under dynamic workloads.
@@ 4.3. Phase-Specific Scheduler
Since Nexus separates prefill and decode into concurrent batches, we can further exploit this with phase-specific optimizations. While SM partitioning governs how much resource each phase gets, our schedulers govern how that resource is consumed. Particularily, we employ scheduler policies tailored to the latency and compute characteristics of each phase.
@@ 4.3.2. Decode Scheduler
The decode phase controls TBT. Unlike prefill, decode scheduling operates at a finer granularity, with each request contributing only a single token to the active batch. Thus, we adopt a simple First-Come-First-Serve (FCFS) policy. FCFS ensures fairness, incurs minimal overhead, and avoids token-level starvation. While more sophisticated schemes that consider context window size or memory bandwidth are theoretically possible, we find that their impact is limited.
@@ 6.1. Experimental Setup
Baselines. We compare Nexus against four representative LLM serving engines, all configured with the same tensor parallelism, chunked prefill, and scheduling budget for fair comparison. We plan to integrate Semi-PD in a future update to provide a more complete intra-GPU PD disaggregation system comparison.
@@ 7. Related Work
Monolithic LLM Serving Systems. Early LLM serving engines focus on maximizing throughput and memory efficiency within a unified execution engine. Orca (Yu et al., 2022) introduces continuous batching to reduce head-of-line blocking. vLLM (Kwon et al., 2023) eliminates KV cache fragmentation via PagedAttention, while SGLang (Zheng et al., 2023) reduces memory usage through RadixAttention. SarathiServe (Agrawal et al., 2024) mixes chunks of prefill and decode requests to better utilize GPU resources. However, all of these designs treat prefill and decode as indistinguishable units within a shared queue. In contrast, Nexus decouples the two phases at the batching level, enabling independent, phase-specific execution and scheduling while preserving compatibility with existing attention mechanisms.
@@
Engine-Level PD Disaggregation Systems. Several systems physically disaggregate prefill and decode across GPUs to better match their resource profiles. Splitwise (Patel et al., 2024) statically assigns phases to different hardware tiers; DistServe (Zhong et al., 2024) improves TTFT/TBT via tiered GPU scheduling; Mooncake (Qin et al., 2025) serves cached KV blocks from storage to reduce compute load; TetriInfer (Hu et al., 2024) routes requests by latency class to isolated replicas. Nexus achieves similar benefits without incurring cross-engine complexity. By performing lightweight intra-engine disaggregation, it enables low-latency execution and efficient KV reuse within a single serving engine.
Intra-GPU PD Disaggregation Systems. Recent systems like Bullet (Lin et al., 2025), Drift (Cui et al., 2025), and semi-PD (Hong et al., 2025) explore intra-GPU disaggregation to reduce cross-device overhead. Bullet builds profile-augmented latency models and uses SLO-driven feedback loops to tune SM partitioning reactively. Drift applies phase-tagged gang scheduling under a contention-free assumption, combining static modeling with latency-aware dispatch. Semi-PD fits inverse-linear latency curves and adjusts SM ratios through runtime feedback control based on latency violations. Nexus takes a proactive approach. First, it introduces a contention-aware analytical model that explicitly captures both diminishing compute returns and dynamic memory bandwidth interference at the operator level. Second, it formulates intra-GPU resource allocation as a dual-objective optimization problem guided by runtime KV-cache usage and solved via fast greedy search. Third, Nexus uses a one-time profiling pass to calibrate per-kernel latency scaling curves, but avoids offline workload tracing or in-deployment feedback fitting, enabling generalization across dynamic traffic and prompt structures.
@@
GPU Multiplexing. Work on GPU multiplexing explores spatial or temporal sharing across models or processes. Systems like GPUlet, Orion, REEF, and Bless (Choi et al., 2022; Strati et al., 2024; Han et al., 2022; Zhang et al., 2025) optimize for small models via fine-grained temporal sharing. MuxServe (Duan et al., 2024) proposes spatial-temporal multiplexing to efficiently serve multiple LLMs, while NanoFlow (Zhu et al., 2024) and Liger (Du et al., 2024) introduce kernel level parallelism that enables concurrent execution of compute bound, network bound, and memory bound operations. Unlike these systems, which focus on generic sharing or inter-model concurrency, Nexus targets intra-GPU, phase-level disaggregation. It dynamically and proactively reallocates GPU resources between prefill and decode phases within a single LLM engine, enabling fine-grained, workload-aware partitioning guided by lightweight performance modeling and coordinated scheduling.
LLM Scheduling Frameworks. Prior works also propose sophisticated schedulers to balance latency and throughput. FastServe (Wu et al., 2023) uses MLFQ with skip-join to avoid prompt variance stalls. VTC (Sheng et al., 2024) and QLM (Patke et al., 2024) target fairness and SLO adherence. Llumnix (Sun et al., 2024) accelerates dynamic scaling via migration-aware scheduling. LightLLM (Gong et al., 2025) and Preble (Srivatsa et al., 2024) improve memory reuse via future prediction or prompt sharing. Yet, all treat each request as a monolithic scheduling unit. Nexus introduces dual schedulers for prefill and decode, each tailored to its phase’s latency sensitivity and compute intensity. This decoupling enables tighter queue control and more efficient GPU resource utilization.